OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

OpenStudy (phi):

can you put it in the form y = mx +b because if you can, then the number in front of the x (the m) is the slope

OpenStudy (anonymous):

I don't know how

OpenStudy (phi):

if you add -5x to both sides of the equation, like this: -5x+5x - 2y= -10 -5x can you simplify this?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

@Samxoxo do you know what's slope?

OpenStudy (phi):

You might want to watch a few videos first, then http://www.khanacademy.org/math/algebra/solving-linear-equations/v/simple-equations

OpenStudy (anonymous):

no!

OpenStudy (ash2326):

Ok, I'll explain you but first one more question Find the value of x? \[x+3=6\]